A 50 gallons tank initially contains 10 gal of fresh water. At t = 0, a brine solution


containing 1 lb of salt per gallon is poured into the tank at the rate of 4 gal/min, while the


well-stirred mixture leaves the tank at the rate of 2 gal/min. Find


a. the amount of time required for overflow to occur


b. the amount of salt in the tank at the moment of overflow

1
Expert's answer
2021-12-02T16:18:53-0500

2.

Let "s(t) =" amount, in lb of salt at time "t." Then we have

"\dfrac{ds}{dt}="(rate of salt into tank) − (rate of salt out of tank)


"\dfrac{ds}{dt}=1(4)-\dfrac{2s}{10+(4-2)t}"

So we get the differential equation


"\dfrac{ds}{dt}+\dfrac{s}{5+t}=4, s(0)=0"


Integrating factor


"\mu=5+t"

"(5+t)\dfrac{ds}{dt}+s=4(5+t)"

"d((5+t)s)=4(5+t)dt"

Integrate


"\int d((5+t)s)=\int4(5+t)dt"

"(5+t)s=20t+2t^2+C"

"s=\dfrac{20t+2t^2+C}{5+t}"

"s(0)=\dfrac{C}{5}=0=>C=0"

"s(t)=\dfrac{20t+2t^2}{5+t}"

a. "V=10+2t=50=>t=20"

The tank overflows at "t=20" min.


b.

"s(20)=\dfrac{20(20)+2(20)^2}{5+20}=48(lb)"